# Sample-Efficient Differentially Private Fine-Tuning via Gradient Matrix Denoising

### v1

@@ Abstract
We address the challenge of sample efficiency in differentially private fine-tuning of large language models (LLMs) using DP-SGD. While DP-SGD provides strong privacy guarantees, the added noise significantly increases the entropy of gradient matrices, disrupting their low-rank structure and slowing optimization. We propose a post-processing algorithm that leverages random matrix theory to denoise gradients, restore low-rank structure, and improve alignment with the original signal. Applied to DP-SGD fine-tuning of RoBERTa on GLUE tasks, our method improves sample efficiency compared to state-of-the-art approaches, substantially reducing training time when optimal performance is not required. This work demonstrates that matrix recovery techniques can enhance the utility of private language model training without compromising privacy guarantees.
@@ 1 Introduction
Applying DP-SGD to language models, while successful, has many challenges. Training large language models (LLMs) with DP-SGD is computationally expensive (Li et al., 2021). Using parameter efficient fine-tuning methods, this challenge has been addressed (Yu et al., 2021). Still, computational cost is higher than the non-private training, because of lower sample efficiency.
@@
The singular values of the gradient matrix undergo a “phase transition” (Baik et al., 2005) when noise is added. If the underlying signal is weak, the singular values of the noisy matrix become indistinguishable from those of pure noise. Figure 1 illustrates this by comparing the sorted singular values of a RoBERTa layer’s gradient matrix before and after DP-SGD noise is applied. In this weak-signal regime, the noisy gradient’s singular values closely follow the “bulk” distribution predicted by the Marchenko–Pastur law (Marčenko & Pastur, 1967; Tao, 2012), making them essentially indistinguishable from pure noise. Thus, when a low-rank signal is too small relative to the noise, it is hidden in the noise and cannot be detected or recovered by examining the singular values and vectors alone. This highlights a fundamental limitation: sufficiently weak signals are undetectable in the presence of strong noise.
However, if some singular values exceed this threshold, the largest singular values of the noisy matrix deviate from the bulk, as shown in Figure 2. This phenomenon is known as the Baik–Ben Arous–Péché (BBP) phase transition (Baik et al., 2005). The extent of these deviations, as well as the alignment between the singular vectors of the noisy and original matrices, can be predicted mathematically (Baik & Silverstein, 2006; Benaych-Georges & Nadakuditi, 2012). These properties enable partial recovery of the original matrix from its noisy observation (Shabalin & Nobel, 2013; Gavish & Donoho, 2014).
@@ 2.1.2 DP-SGD
In DP-SGD, each individual gradient is first clipped so that its norm does not exceed the threshold $C$. The clipped gradients are then summed, and Gaussian noise with entries drawn from $\mathcal{N}(0,\sigma C)$ is added. Finally, the result is averaged over the batch:
@@
$\tilde{\bm{g}}_{\mathcal{B}}$ $=\frac{1}{B}\left(\bar{\bm{g}}_{\mathcal{B}}+\bm{w}\right),\quad\bm{w}_{j}\sim\mathcal{N}(0,\sigma C)$
@@
One other important aspect of DP-SGD is the sampling mechanism. Most of the privacy accountants rely on the notion of ”sampling with replacement” to ensure that each data point is independently and identically distributed (i.i.d.) during training. So the sampling process must account for this to maintain the desired privacy guarantees.
Algorithm 1 DP-SGD
0: Dataset $D$, loss function $f(\theta,x)$, model parameters $\theta$, sampling rate $\rho$, clipping norm $C$, noise multiplier $\sigma$, optimizer $\mathcal{O}$, number of steps $T$
for $t\leftarrow 1$ to $T$ do
  Sample a batch $\mathcal{B}$ from $D$ using Poisson sampling with rate $\rho$
  for each $i\in\mathcal{B}$ do
  end for
[…]
end for
2.1.3 Post processing invariance
A fundamental property of differential privacy is its invariance under post-processing. This means that no adversary, regardless of the method applied to the output of a differentially private algorithm, can reduce its privacy guarantees or extract more information about the original dataset. In other words, post-processing cannot make the output less private, providing strong protection against attempts to compromise privacy. While previous work has leveraged this property to improve the utility of the DP-SGD algorithm (Zhang et al., 2024; Balle & Wang, 2018), none have utilized results from random matrix theory for the post-processing function. To our knowledge, this is the first work to apply such results in the context of DP-SGD.
2.2 Singular value distribution of gradients
2.3 Low rank matrix estimations
Low rank matrix reconstruction is a rich sub-field of signal processing (Donoho et al., 2018; Gavish & Donoho, 2014; Shabalin & Nobel, 2013). Assuming the rank of the signal matrix $\bm{X}\in\mathbb{R}^{m\times n}$ is $k$, we can use the SVD decomposition to write it as
$\bm{X}=\sum_{i=1}^{k}\lambda_{i}\bm{u}_{i}\bm{v}_{i}^{T}$
With the mentioned notation we have
This is an increase in the value of the singular value, which is usual in random matrix theory. It is important to note that these results are typically stated in the asymptotic regime, where the matrix dimensions grow to infinity and the noise variance may scale with the dimensions, often under specific assumptions on the ratio $m/n$. In practical, finite-dimensional settings, these approximations may incur some error. The precise rate of this error in finite dimensions is not addressed here and could be an interesting direction for further study. The derivation of these results from their asymptotic forms is postponed to the appendix A.
Also, assuming all of the eignvalues of $\bm{X}$ are distinct, and if $\lambda_{i}>\sigma\sqrt[4]{mn}$, following Lemma 3 of Gavish & Donoho (2014) or proposition 9 of Shabalin & Nobel (2013), we can write
and
However if $\lambda_{i}\leq\sigma\sqrt[4]{mn}$, then
2.3.2 Matrix Denoising
Matrix denoising methods aim to recover the underlying signal matrix $\bm{X}$ from its noisy observation $\tilde{\bm{X}}$ by leveraging the low-rank structure of the signal. Many of these methods shrink the singular values of the noisy matrix. One such approach is the so-called optimal method discussed in Shabalin & Nobel (2013); Donoho et al. (2018), which outputs a low-rank matrix.
Optimal Denoising
where, following Shabalin & Nobel (2013), the optimal coefficients are
3 Methodology
In this section, we introduce our post-processing method, which leverages equations 7 and 8 to denoise the gradients produced by DP-SGD before they are passed to the optimizer. Note that some of the performance loss in DP-SGD, compared to non-private training, is due to gradient clipping (Bu et al., 2023). However, this work does not address the performance drop from clipping; we focus exclusively on mitigating the loss caused by the added noise.
3.1 Framework
For tracking this value, we define the Improvement at step $t$ as
If we can come up with such a denoising function, we hope to improve the sample efficiency of DP-SGD by making the noisy gradients more closely resemble the true (clipped) gradients. Having such a denoising function, we can change the DP-SGD algorithm as follows:
Algorithm 2 DP-SGD with Denoising
@@ 3.1 Framework
We expect that if the improvement at each step $t$ is consistently non negative, $\mathrm{Improvement}(t)\geq 0$, then the denoising function is effectively aligning the noisy gradients with the true (clipped) gradients, leading to faster convergence of the DP-SGD algorithm. The following sections will detail the implementation of the denoising function which are mainly based on the results reviewed in Section 2.3.2.
@@ 3.2 Denoising Function
Where if a layer $\bm{W}_{i}$ is not a linear layer, we simply set $\mathrm{Denoise}(\tilde{\bm{g}}_{W_{i}})=\tilde{\bm{g}}_{W_{i}}$.
For linear layers, we modify the so called “optimal” denoising method in two ways
  • Only applying gradient if the singular values of the noisy layer gradient are larger than a preset multiple of $\sigma(\sqrt{n}+\sqrt{m})$, the largest singular value of the bulk.
@@
3.2.1 Why threshold is needed?
It is important to recognize that the results in Section 2.3.2 are derived in asymptotic settings. For instance, the theory predicts that if all singular values of the signal matrix are less than $\sigma\sqrt[4]{nm}$, or equivalently, if all singular values of the noisy matrix are less than $\sigma(\sqrt{n}+\sqrt{m})$, then the inner products between the left (or right) singular vectors of the signal and noisy matrices should be zero. In that case, the optimal denosing algorithm returns the zero matrix as the optimum result and states that it is the best one can get. However, in practice and for finite-dimensional matrices, this is not true, and the noisy gradient, even if its singular value are small, usually still has some positive cosine similarity with the original gradient. As a result, the denoising algorithm does not always improve the alignment between the noisy and clipped gradients.
Fortunately, we identified a simple heuristic criterion to guide when the denoising algorithm should be applied. Specifically, we require that the largest singular value of the noisy gradient matrix exceeds the threshold $\kappa\,\sigma(\sqrt{n}+\sqrt{m})$ before applying the denoising algorithm. Our observations show that when this condition is met, denoising tends to improve the alignment; otherwise, it may decrease it. For choosing the value of $\kappa$, we tuned it on the SST dataset while training the roberta-base model by choosing the best value from the set $\{1.01,1.02,1.05,1.1\}$, and used the same value for all the other model/datset pairs.
@@ 3.2.2 Why norm correction is needed?
Norm correction is needed because of the possible big reduction in the norm of the denoised gradient compared to the noisy gradient. This may in turn cause the global improvement at time $t$ to become negative. Why this might happen, and how rescaling helps prevent it, will be explained in the appendix B.3, as well as results of an abalation study on the effect of norm correction.
@@ 4 Experiments
We keep every other hyper-parameter the same as Yu et al. (2021), including batch size, learning rate, weight decay, and clipping norm. Looking at the tables 1 and 2, we can see that our method consistently improves the sample efficiency of DP-SGD across all datasets and model sizes. Improvements range from $20\%$ to $100\%$ in the number of steps required to reach $90\%$ and $95\%$ of the SOTA performance. Also, we achieved higher performance in five out of eight cases for the final accuracy at $400$ steps.
@@
To explain why this method converges faster, we can check the improvement in cosine similarity between the denoised and noisy gradients with respect to the clipped gradients. As shown in Figure 4, the denoising method consistently improves the alignment between the noisy and clipped gradients throughout the training process, which aligns with our explanation of why the speedup happens.

### v2

@@ Abstract
We address the challenge of sample efficiency in differentially private fine-tuning of large language models (LLMs) using DP-SGD. While DP-SGD provides strong privacy guarantees, the added noise significantly increases the entropy of gradient matrices, disrupting their low-rank structure and slowing optimization. We propose a post-processing algorithm that leverages random matrix theory to denoise gradients, restore low-rank structure, and improve alignment with the original signal. Applied to DP-SGD fine-tuning of RoBERTa model family on GLUE tasks and Qwen and Llama family on DART and E2E datasets, our method improves sample efficiency compared to state-of-the-art approaches, substantially reducing training time when optimal performance is not required. This work demonstrates that matrix recovery techniques can enhance the utility of private language model training without compromising privacy guarantees.
@@ 1 Introduction
Applying DP-SGD to language models, while successful, has many challenges. Training large language models (LLMs) with DP-SGD is computationally expensive (Li et al., 2021). Using parameter efficient fine-tuning methods, this challenge has been addressed (Yu et al., 2021). Still, computational cost is higher than the non-private training, because of lower sample efficiency. This can be viwed, for example, in Figure 1.
@@
The singular values of the gradient matrix undergo a “phase transition” (Baik et al., 2005) when noise is added. If the underlying signal is weak, the singular values of the noisy matrix become indistinguishable from those of pure noise. Figure 4 illustrates this by comparing the sorted singular values of a RoBERTa layer’s gradient matrix before and after DP-SGD noise is applied. In this weak-signal regime, the noisy gradient’s singular values closely follow the “bulk” distribution predicted by the Marchenko–Pastur law (Marčenko and Pastur, 1967; Tao, 2012), making them essentially indistinguishable from pure noise. Thus, when a low-rank signal is too small relative to the noise, it is hidden in the noise and cannot be detected or recovered by examining the singular values and vectors alone. This highlights a fundamental limitation: sufficiently weak signals are undetectable in the presence of strong noise.
However, if some singular values exceed this threshold, the largest singular values of the noisy matrix deviate from the bulk, as shown in Figure 5. This phenomenon is known as the Baik–Ben Arous–Péché (BBP) phase transition (Baik et al., 2005). The extent of these deviations, as well as the alignment between the singular vectors of the noisy and original matrices, can be predicted mathematically (Baik and Silverstein, 2006; Benaych-Georges and Nadakuditi, 2012). These properties enable partial recovery of the original matrix from its noisy observation (Shabalin and Nobel, 2013; Gavish and Donoho, 2014).
@@ 2.1.2 DP-SGD
In DP-SGD, each individual gradient is first clipped so that its norm does not exceed the threshold $C$. The clipped gradients are then summed, and Gaussian noise with entries drawn from $\mathcal{N}(0,\sigma^{2}C^{2})$ is added. Finally, the result is averaged over the batch:
@@
$\tilde{\bm{g}}_{\mathcal{B}}$ $=\frac{1}{B}\left(\bar{\bm{g}}_{\mathcal{B}}+\bm{w}\right),\quad\bm{w}_{j}\sim\mathcal{N}(0,\sigma^{2}C^{2})$
@@
Algorithm 1 DP-SGD (with Denoising)
@@
2.1.3 Post processing invariance
A fundamental property of differential privacy is its invariance under post-processing. This means that no adversary, regardless of the method applied to the output of a differentially private algorithm, can reduce its privacy guarantees or extract more information about the original dataset. In other words, post-processing cannot make the output less private, providing strong protection against attempts to compromise privacy. While previous work has leveraged this property to improve the utility of the DP-SGD algorithm (Zhang et al., 2024; Balle and Wang, 2018), none have utilized results from random matrix theory for the post-processing function. To our knowledge, this is the first work to apply such results in the context of DP-SGD.
2.2 Singular value distribution of gradients
2.3 Low rank matrix estimations
Low rank matrix reconstruction is a rich sub-field of signal processing (Donoho et al., 2018; Gavish and Donoho, 2014; Shabalin and Nobel, 2013). Assuming the rank of the signal matrix $\bm{X}\in\mathbb{R}^{m\times n}$ is $k$, we can use the SVD decomposition to write it as
$\bm{X}=\sum_{i=1}^{k}\lambda_{i}\bm{u}_{i}\bm{v}_{i}^{T}$
With the mentioned notation we have
This is an increase in the value of the singular value, which is usual in random matrix theory. It is important to note that these results are typically stated in the asymptotic regime, where the matrix dimensions grow to infinity and the noise variance may scale with the dimensions, often under specific assumptions on the ratio $m/n$. In practical, finite-dimensional settings, these approximations may incur some error. The precise rate of this error in finite dimensions is not addressed here and could be an interesting direction for further study. The derivation of these results from their asymptotic forms is postponed to the appendix A.2.
Also, assuming all of the eignvalues of $\bm{X}$ are distinct, and if $\lambda_{i}>\sigma\sqrt[4]{mn}$, following Lemma 3 of Gavish and Donoho (2014) or proposition 9 of Shabalin and Nobel (2013), we can write
and
However if $\lambda_{i}\leq\sigma\sqrt[4]{mn}$, then
2.3.2 Matrix Denoising
Matrix denoising methods aim to recover the underlying signal matrix $\bm{X}$ from its noisy observation $\tilde{\bm{X}}$ by leveraging the low-rank structure of the signal. Many of these methods shrink the singular values of the noisy matrix. One such approach is the so-called optimal method discussed in Shabalin and Nobel (2013); Donoho et al. (2018), which outputs a low-rank matrix.
Optimal Denoising
where, following Shabalin and Nobel (2013), the optimal coefficients are
Computational Overhead
Although the additional computation required for matrix denoising may seem significant, with careful implmenetation and also utilizing the parallelizable nature of the optimal denoising algorithm, the overhead can be kept minimal. We draw readers attention to three properties that can be leveraged to reduce the computational overhead. 1) First, the fact that in equation 7, the $F^{-1}_{\sigma,n,m}(\tilde{\lambda}_{i})$ can be computed independently for each singular value, and the same applies to the computation of $\eta_{i}$. 2) Second, the SVD computation, which is the most computationally expensive part of the algorithm, can be computed in a batched manner for all the layers of the neural network with similar dimensions. And 3) Third, by utilizing approximate methods when acceptable, the SVD computation can be further accelerated. While if implmeneted naïvely, the overhead can be just short of $3\%$ of the total training time, combining the mentioned strategies, the computational overhead of the denoising step can be reduced to less than $1\%$ of the total training time in our experiments (Table 1). The efficient implementation is available as part of the supplementary material.
Table 1: Training time in seconds for different methods of DP-SGD fine-tuning of RoBERTa on SST-2 dataset with the setting of section 4. The overhead is computed as the percentage increase in training time compared to regular DP-SGD.
Model Method Train Time Overhead
RoBERTa Base Regular DP-SGD 1846
Naïve denoising implementation 1899 2.87%
Efficient denoising implementation 1861 0.81%
RoBERTa Large Regular DP-SGD 5658
Naïve denoising implementation 5781 2.17%
Efficient denoising implementation 5700 0.74%
3 Methodology
In this section, we introduce our post-processing method, which leverages equations 7 and 8 to denoise the gradients produced by DP-SGD before they are passed to the optimizer. First, we establish that the slower convergence observed with DP-SGD (compared to non-private training) is primarily caused by the added Gaussian noise, not by gradient clipping. Although prior work often attributes the loss gap between DP-SGD and non-private training to clipping (Bu et al., 2023), these are distinct phenomena. Figure 1 shows that the final validation loss is similar for DP-SGD with a zero noise multiplier ($\sigma=0$) and for DP-SGD using the noise level required for privacy; however, convergence toward that final loss is significantly slower when noise is added. This indicates that the added noise is the main driver of DP-SGD’s slower convergence. This is why we focus on denoising the gradients in our method. More similar experiments comparing clipping alone versus DP-SGD can be found in appendix B.
Figure 1: Comparison of validation loss between DP-SGD and training using per-sample clipping only (no noise injected). This is for training RoBERTa base model on sst-2 dataset. The Setting are similar to what is described in 4, except that for only clipping method, we do not add the gaussian noise and only perform the per-sample clipping. It is evident that the slowdown in convergence is more pronounced for DP-SGD than for clipping alone.
3.1 Framework
For tracking this value for evaluation purposes, we define the Improvement at step $t$ as
If we can come up with such a denoising function, we hope to improve the sample efficiency of DP-SGD by making the noisy gradients more closely resemble the true (clipped) gradients. Having such a denoising function, we can change the DP-SGD algorithm as in Algorithm LABEL:alg:dp_sgd_denoised.
We expect that if the improvement at each step $t$ is consistently non negative, $\mathrm{Improvement}(t)\geq 0$, then the denoising function is effectively aligning the noisy gradients with the true (clipped) gradients, leading to faster convergence of the DP-SGD algorithm. It is important to note that the improvement function is only used for evaluation porpuses and is not used for the algorithm itself, as doing so would violate the differential privacy criteria. The following sections will detail the implementation of the denoising function which are mainly based on the results reviewed in Section 2.3.2. The guiding principle behind adapting the matrix denoising methods to our task is that the denoiser should increase alignment with the (private) clipped gradient, while itself using no additional private information. This principle helps us in two design choices. One gives us a way to adapt the asymptotic formulas to finite dimension, and the other helps us to generalize the denoising algorithm from operating on a single matrix to operating on a collection of matrices (the layers of the neural network).
@@ 3.2 Denoising Function
Where if a layer $\bm{W}_{i}$ is not a linear layer, we simply set $\mathrm{Denoise}(\tilde{\bm{g}}_{W_{i}})=\tilde{\bm{g}}_{W_{i}}$. With this notation, we can also define the per-layer improvement as
$\mathrm{Improvement}_{i}(t)=\mathrm{cos}(\mathrm{Denoise}(\tilde{\bm{g}}_{t}\!\mid_{\bm{W}_{i}}),\bar{\bm{g}}_{t}\!\mid_{\bm{W}_{i}})-\mathrm{cos}(\tilde{\bm{g}}_{t}\!\mid_{\bm{W}_{i}},\bar{\bm{g}}_{t}\!\mid_{\bm{W}_{i}}).$
For linear layers, we modify the so called “optimal” denoising method in two ways.
  • To adapt the assymptotic formulas to finite dimension we only apply optimal denoising if the singular values of the noisy layer gradient are larger than a preset multiple of $\sigma(\sqrt{n}+\sqrt{m})$, where $n,m$ are dimensions of that linear layer. When the singular value is larger than the required threshold, we apply the optimal denoising function.
@@
3.2.1 Why threshold is needed and why this specific value?
It is important to recognize that the results in Section 2.3.2 are derived in asymptotic settings. For instance, the theory predicts that if all singular values of the signal matrix are less than $\sigma\sqrt[4]{nm}$, or equivalently, if all singular values of the noisy matrix are less than $\sigma(\sqrt{n}+\sqrt{m})$, then the inner products between the left (or right) singular vectors of the signal and noisy matrices should be zero. In that case, the optimal denosing algorithm returns the zero matrix as the optimum result and states that it is the best one can get. However, in practice and for finite-dimensional matrices, this is not true, and the noisy gradient, even if its singular value are small, usually still has some positive cosine similarity with the original gradient. As a result, the denoising algorithm does not always improve the alignment between the noisy and per-layer clipped gradients.
Fortunately, we identified a simple RMT-based criterion to decide when to apply the denoiser. Concretely, we only denoise a layer if the largest singular value of its noisy gradient exceeds $\kappa\,\sigma(\sqrt{n}+\sqrt{m})$. Choosing $\kappa=1$ prevents denoising in cases where the optimal estimator would return the zero matrix. This threshold is motivated by the phase transition in equation 3: $\sigma(\sqrt{n}+\sqrt{m})$ is the maximal singular value of a pure-noise matrix (signal equal to zero). Thus, RMT indicates when denoising can meaningfully improve alignment; if the estimator would output the zero matrix, the theoretical prediction is that there is no signal to recover.
Our observations show that for correct value of $\kappa$, denoising tends to improve the alignment, otherwise it may reduce the alignment . For choosing the value of $\kappa$, we tuned it on the SST dataset while training the RoBERTa-base model by choosing the best value from the set $\{1.01,1.02,1.05,1.1\}$, and used the same value for all the other model/datset pairs. We choose this set to have values greater than $1$ to avoid the cases where denoising outputs zero matrix. Also, we want to keep the values as close to $1$ as possible to have more layers denoised. The trade-off here is to have a big enough $\kappa$ so the denoising improves the alignment, and to have it small enough so that we get enough per-layer gradients denoised to get the most out of the alignment improvement. The best value we found was $\kappa=1.02$ (Figure 2). We also like to emphasis that while we could have tuned $\kappa$ for each model/dataset pair, we refrained from doing so to show the robustness of our method to this hyperparameter. The value $\kappa=1.02$ worked well accross all the model/dataset pairs we tried.
Figure 2: Effect of different values of $\kappa$ on the per-layer improvements when training RoBERTa base model on sst-2 dataset. The Setting are similar to what is described in 4. It is evident that $\kappa=1.02$ is the smallest $\kappa$ which has dominantly positive per-layer improvement.
@@ 3.2.2 Why norm correction is needed?
The scope in which the denoising function from the RMT works is to improve the alignment between the noisy and clipped per-layer gradients. There is no extension of the RMT method to a combination of different layers, and the relative scaling between them. For generalizing to a method for improving alignment of the whole gradient vector, we are going to use the following theorem, which states that if we improve the alignment of each component of a vector, and keep their norms the same, then the overall alignment will also improve.
Theorem 1.
Let $x=(x_{1},\dots,x_{c})\in\mathbb{R}^{n}$ be a target vector, with $x_{i}\in\mathbb{R}^{n_{i}}$ and $\sum_{i=1}^{c}n_{i}=n$. Let $y=(y_{1},\dots,y_{c}),y^{\prime}=(y^{\prime}_{1},\dots,y^{\prime}_{c})\in\mathbb{R}^{n}$ be estimations of $x$, with $y_{i},y^{\prime}_{i}\in\mathbb{R}^{n_{i}}$. If we have
  1. (i)
    $\mathrm{cos}(y_{i},x_{i})\leq\mathrm{cos}(y^{\prime}_{i},x_{i})$, or all $i\in\{1,\dots,c\}$. and
  2. (ii)
    $\|y_{i}\|=\|y^{\prime}_{i}\|$, for all $i\in\{1,\dots,c\}$.
Then, we have $\mathrm{cos}(y,x)\leq\mathrm{cos}(y^{\prime},x)$.
[…]
On another note, the assumption (ii) in theorem 1 is the reason we need to do the norm correction in equation 9. Without this correction, even if the per-layer alignment improves, there is no guarantee that the overall alignment will also improve. To see the effect of the norm correction, we have done abalation studies for the effect of norm correction. The results for the case of training RoBERTa base model on SST-2 dataset are presented in figure 3. It is evident that without the norm correction, the improvement is not consistent, and even negative. This shows the importance of the norm correction step in our denoising function. More abalation results can be found in the appendix D.
Figure 3: Comparison of whole gradient improvement when using norm correction and not using it. This is for training RoBERTa base model on sst-2 dataset. The Setting are similar to what is described in 4. It is evident that using norm correction results in more consistent improvement.
Also, on the last note, we like to note that the norm of the denosied gradient in this method is exactly the nosie of the noisy gradient. This will prevent any possible issues with convergence of the optimizer due to unexpected changes in the gradient norm.
@@ 4 Experiments
We keep every other hyper-parameter the same as Yu et al. (2021), including batch size, learning rate, weight decay, and clipping norm. Looking at the tables 2 and 3, we can see that our method consistently improves the sample efficiency of DP-SGD across all datasets and model sizes. Improvements range from $\mathbf{20\%}$ to $\mathbf{100\%}$ in the number of steps required to reach $90\%$ and $95\%$ of the SOTA performance. Also, we achieved higher performance in five out of eight cases for the final accuracy at $400$ steps.
@@
We also report additional language-generation experiments on the E2E benchmark (Novikova et al., 2017) and DART (Nan et al., 2021) using Qwen3 and Llama3.2 models (Team, 2025; Grattafiori et al., 2024) (see Appendix E for hyperparameters and more details). After 50 training steps (Table 4), our denoising method outperforms the baseline in 49 of 50 model/dataset/metric combinations, demonstrating substantial sample-efficiency gains under tight iteration budgets. These experiments use models of up to 4B parameters, indicating the approach scales to larger models.